Assign an oxidation number to each element in the reaction. CaF(s)+H2SO4(aq)=CaSO4(aq)+2HF(g)
In CaF2 the oxidation number of Ca is?
In H2SO4 the oxidation number of H is and that of S is?
In CaSO4 the oxidation number of Ca is and that of S is?
In HF the oxidation number of H is?
1
Expert's answer
2020-05-04T01:52:50-0400

Solution.

"{x\atop Ca}{-1\atop F_2}; x+(-1)\sdot2=0; x=+2; {+2\atop Ca}{-1\atop F_2};"

"{+1\atop H_2}{x\atop S}{-2\atop O_4}; 2\sdot(+1)+x+(-2)\sdot4=0; x=+6;{+1\atop H_2}{+6\atop S}{-2\atop O_4};"


"{x\atop Ca}{+6\atop S}{-2\atop O_4}; x+(+6)+(-2)\sdot4=0;x=+2; {+2\atop Ca}{+6\atop S}{-2\atop O_4};"


"{x\atop H}{-1\atop F}; x+(-1)=0;x=+1;{+1\atop H}{-1\atop F};"


"{+2\atop Ca}{-1\atop F_2}+{+1\atop H_2}{+6\atop S}{-2\atop O_4}={+2\atop Ca}{+6\atop S}{-2\atop O_4}+2{+1\atop H}{-1\atop F}" ;


Answer: "{+2\atop Ca}{-1\atop F_2}+{+1\atop H_2}{+6\atop S}{-2\atop O_4}={+2\atop Ca}{+6\atop S}{-2\atop O_4}+2{+1\atop H}{-1\atop F}" .